Problem: NEO 2000e and NEO 4000e units containing HP LTO-3 FC tape drives do not respond via the fibre channel port. The NEO E-series Touchscreen and Web Management Interface show the drives status correctly.
Solution: Remove the back cover of the drive shoe and move the pigtail connection from the left side fibre channel port (Port B) to the right side fibre channel port (Port A) then reassemble the drive shoe and try again. Refer to the attachment to help identify the correct fibre channel port (Port A).
Note: This problem has been fixed in manufacturing. This Knowledge center document was written to cover any units that may have already been shipped before the change.